# Simulations of stochastic fluid dynamics near a critical point in the phase diagram

### v1

@@ III Numerical method
$\phi^{\rm trial}(\vec{x},t+\Delta t)$ $=\phi(\vec{x},t)+q_{\mu}\,,$ (7)
$\phi^{\rm trial}(\vec{x}+\hat{\mu},t+\Delta t)$ $=\phi(\vec{x}+\hat{\mu},t)-q_{\mu}\,,$ (8)
$q_{\mu}$ $=\sqrt{2\Gamma T(\Delta t)}\,\zeta\,,$ (9)
where $\zeta$ is a Gaussian random variable with zero mean and unit variance, and $\hat{\mu}$ is an elementary lattice vector in the direction $\mu=1,\ldots,d$. The update is accepted with probability ${\rm min}(1,e^{-\Delta{\cal H}/T})$. This algorithm is based on the observation that the average update $\langle[\phi(\vec{x},t+\Delta t)-\phi(\vec{x},t)]\rangle$ realizes the diffusion equation, and the second moment $\langle[\phi(\vec{x},t+\Delta t)-\phi(\vec{x},t)]^{2}\rangle$ reproduces the noise term, see Florio et al. (2022); Chattopadhyay et al. (2023).
@@
$\pi_{\nu}^{\rm trial}(\vec{x},t+\Delta t)$ $=\pi^{T}_{\nu}(\vec{x},t)+r_{\nu\mu}\,,$ (10)
$\pi_{\nu}^{\rm trial}(\vec{x}+\hat{\mu},t+\Delta t)$ $=\pi^{T}_{\nu}(\vec{x}+\hat{\mu},t)-r_{\nu\mu}\,,$ (11)
$r_{\nu\mu}$ $=\sqrt{2\eta T(\Delta t)}\,\xi_{\nu}\,,$ (12)
where $r_{\mu\nu}$ is a random flux and $\xi_{\nu}$ are Gaussian random variables with $\langle\xi_{\mu}\xi_{\nu}\rangle=\delta_{\mu\nu}$. Again, the update is accepted with probability ${\rm min}(1,e^{-\Delta{\cal H}/T})$. After a sweep through the lattice we project on the transverse component of the momentum density, $\pi_{\mu}^{T}(\vec{x},t)=P^{T}_{\mu\nu}\pi_{\nu}(\vec{x})$. The projection is carried out in Fourier space.
@@
In the continuum limit, these equations conserve the integrals of $\phi$ and $\pi_{\mu}^{T}$, as well as the Hamiltonian ${\cal H}$. We have found that it is important to preserve these conservation laws in the lattice theory to the greatest extent possible. Using the skew discretized derivatives introduced by Morinishi et al. Morinishi et al. (1998) it is possible to construct an advection step that conserves integrals of the kinetic energy terms $\frac{1}{2\rho}(\pi^{T}_{\mu})^{2}+\frac{1}{2}(\nabla_{\mu}\phi)^{2}$ exactly Chattopadhyay et al. (2024). We integrate the equations (13,14) using the strongly stable third-order Runge-Kutta scheme of Shu and Osher Shu and Osher (1988). The fields $\phi$ and $\pi_{\mu}^{T}$ satisfy conservation laws after projection, and the total energy is conserved to very good accuracy.
@@ IV Results
Figure 1: Renormalized viscosity as a function of the bare viscosity in four different models: 1) pure momentum diffusion (no mode-couplings), 2) self-advection ($\pi^{T}$ only couples to itself), 3) critical model H0 (mutual advection of $\phi$ and $\vec{\pi}_{T}$ only), and 4) model H at the critical point. All data were taken in lattice volume $V=L^{3}$ with $L=48$ and with $\rho=1$.
@@
Figure 2: Dynamic order parameter correlation function $C(t,\vec{k})$ for the second non-trivial momentum mode and for different values of $L$ plotted as a function of the scaled time variable. This figure shows data taken at $\eta_{0}=10^{-2}$ for $L=40$ and $48$. Data collapse occurs for $z\simeq 3.01$, and the model B value $z=4$ is clearly excluded.
@@
Figure 3: Scaling exponent $z$ extracted from the dynamic correlation functions for different values of the bare viscosity $\eta_{0}$ in model H0. We determined $z$ by comparing the correlation function for two different volumes, $L=40$ and $48$. We also show the prediction of the Kawasaki approximation, equ. (17); the error band is defined by varying the correlation length in the range $\xi\in[L/2\pi,L/2]$.
The values $z=3$ and $4$ are approximate, a more sophisticated calculation using the $\epsilon$-expansion gives $z=3.07$ and $z=3.96$, but the Kawasaki function is a very good approximation to the behavior of real fluids Swinney and Henry (1973). We conclude from Eq. (17) that in any finite volume, we are likely to observe a scaling exponent between 3 and 4, and that observing the model H scaling exponent requires a combination of a very large correlation length and a small viscosity.
Since the renormalized viscosity is smaller in model H0 (see Fig. 1) we have explored the scaling behavior in model H0. At the critical point $m^{2}=m_{c}^{2}$ we compute the correlation for a range of values of $\eta$ and $L^{3}$. For any given $\eta$ we look for data collapse when comparing different $L$ in order to determine the value of $z$. This is shown in Fig. 2 for $\eta=0.01$ and $L=40$ and $L=48$. We observe that scaling works very well. We then plot the extracted value of $z$ as a function of $\eta$. The result is shown in Fig. 3, and compared to the Kawasaki prediction. We observe the expected crossover from $z\simeq 4$ at large viscosity to $z\simeq 3$ for small viscosity. For $\eta=10^{-2}$ we obtain the dynamical exponent $z\simeq 3.013\pm 0.058$, consistent with the prediction of the two-loop $\epsilon$-expansion, $z\simeq 3.0712$ Adzhemyan et al. (1999).
@@
We consider model H with parameters relevant to a QCD critical endpoint. We take $a=0.75$ fm and $\Delta t=0.3$ fm, and consider a temperature $T=130$ MeV. Then, the enthalpy density of a non-interacting quark-gluon plasma corresponds to $\rho=11.1$ in lattice units. A viscosity to entropy density ratio $\eta/s=1/(4\pi)$ implies $\eta=0.50$ 444 Note that $\eta$ is somewhat bigger than the minimum value of $\eta_{R}$ shown in Fig. 1. This is consistent with the observation by Kovtun et al. Kovtun et al. (2011) that fluctuations lead to a bound on $\eta$ which is slightly weaker than the string theory bound $\eta/s\geq 1/(4\pi)$.. We take a simulation volume $V_{0}=L^{3}$ and measure the order parameter in half the simulation volume. The results are shown in Fig. 4. The correlation functions satisfy dynamical scaling for all values of $n$, but the relaxation time depends on $n$. In particular, $G_{2}(t)$ decays more quickly than $G_{1}(t)$, but the relaxation rate of $G_{3}(t)$ is intermediate between $G_{1}(t)$ and $G_{2}(t)$. These results are not compatible with simple mean field models. Note that at the critical point the correlation length is only limited by $L$, and equilibration is extremely slow: In physical units, the $1/e$ decay time in Fig. 4 exceeds $10^{3}$ fm.
@@ V Summary and outlook:
There are a number of important future directions that we wish to pursue. One important goal is to study realistic backgrounds, corresponding to an expanding and cooling fluid. There are two basic approaches that one might follow. One approach is to extend the methods described in this work to stochastic relativistic fluid dynamics, where we retain all the degrees of freedom of the fluid (both shear and sound modes). This could be accomplished along the lines recently proposed in Basar et al. (2024). Another option is to couple the model described here to a deterministic background flow, obtained from conventional fluid dynamic simulations.
Acknowledgments: This work is supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics through the Contracts DE-FG02-03ER41260 (T.S.) and DE-SC0020081 (V.S.). We used computing resources provided by the North Carolina State University High Performance Computing Services Core Facility (RRID:SCR-022168), as well as resources funded by the Wesley O. Doggett endowment. We thank Andrew Petersen for assistance in working with the NCSU HPC infrastructure.

### v2

@@ III Numerical method
$\phi^{\rm trial}(\vec{x},t+\Delta t)$ $=\phi(\vec{x},t)+q^{(\mu)}\,,$ (7)
$\phi^{\rm trial}(\vec{x}+\hat{\mu},t+\Delta t)$ $=\phi(\vec{x}+\hat{\mu},t)-q^{(\mu)}\,,$ (8)
$q^{(\mu)}$ $=\sqrt{2\Gamma T(\Delta t)}\,\zeta^{(\mu)}\,,$ (9)
where each $\zeta^{(\mu)}$ is a Gaussian random variable with zero mean and unit variance, and $\hat{\mu}$ is an elementary lattice vector in the direction $\mu=1,\ldots,d$. The update is accepted with probability ${\rm min}(1,e^{-\Delta{\cal H}/T})$. This algorithm is based on the observation that the average update $\langle[\phi(\vec{x},t+\Delta t)-\phi(\vec{x},t)]\rangle$ realizes the diffusion equation, and the second moment $\langle[\phi(\vec{x},t+\Delta t)-\phi(\vec{x},t)]^{2}\rangle$ reproduces the noise term, see Florio et al. (2022); Chattopadhyay et al. (2023).
@@
$\pi_{\nu}^{\rm trial}(\vec{x},t+\Delta t)$ $=\pi^{T}_{\nu}(\vec{x},t)+r_{\nu}^{(\mu)}\,,$ (10)
$\pi_{\nu}^{\rm trial}(\vec{x}+\hat{\mu},t+\Delta t)$ $=\pi^{T}_{\nu}(\vec{x}+\hat{\mu},t)-r_{\nu}^{(\mu)}\,,$ (11)
$r_{\nu}^{(\mu)}$ $=\sqrt{2\eta T(\Delta t)}\,\xi_{\nu}^{(\mu)}\,,$ (12)
where $r_{\nu}^{(\mu)}$ is a random flux and $\xi_{\nu}^{(\mu)}$ are Gaussian random variables with $\langle\xi_{\mu}^{(\alpha)}\xi_{\nu}^{(\beta)}\rangle=\delta_{\mu\nu}\delta^{\alpha\beta}$. Again, the update is accepted with probability ${\rm min}(1,e^{-\Delta{\cal H}/T})$. After a sweep through the lattice we project on the transverse component of the momentum density, $\pi_{\mu}^{T}(\vec{x},t)=P^{T}_{\mu\nu}\pi_{\nu}(\vec{x})$. The projection is carried out in Fourier space.
@@
In the continuum limit, these equations conserve the integrals of $\phi$ and $\pi_{\mu}^{T}$, as well as the Hamiltonian ${\cal H}$. We have found that it is important to preserve these conservation laws in the lattice theory to the greatest extent possible. Using the skew discretized derivatives introduced by Morinishi et al. Morinishi et al. (1998) it is possible to construct an advection step that conserves integrals of the kinetic energy terms $\frac{1}{2\rho}(\pi^{T}_{\mu})^{2}+\frac{1}{2}(\nabla_{\mu}\phi)^{2}$ exactly Chattopadhyay et al. (2024). We integrate the equations (13,14) using the strongly stable third-order Runge-Kutta scheme of Shu and Osher Shu and Osher (1988). The fields $\phi$ and $\pi_{\mu}^{T}$ satisfy conservation laws after projection, and the total energy is conserved to very good accuracy. This statement can be quantified in terms of the observed shift in the critical mass (see Sect. IV), which is about 1%.
@@ IV Results
Figure 1: Renormalized viscosity as a function of the bare viscosity in four different models: 1) pure momentum diffusion (no mode-couplings), 2) self-advection ($\pi^{T}$ only couples to itself), 3) critical model H0 (mutual advection of $\phi$ and $\vec{\pi}_{T}$ only), and 4) model H at the critical point. All data were taken in lattice volume $V=L^{3}$ with $L=48$ and with $\rho=1$. Model H results were offset horizontally for better visibility.
@@
Figure 2: Dynamic order parameter correlation function $C(t,\vec{k})$ for the second non-trivial momentum mode and for different values of $L$ plotted as a function of the scaled time variable. This figure shows data taken at $\eta=10^{-2}$ for $L=40$ and $48$. Data collapse occurs for $z\simeq 3.01$, and the model B value $z=4$ is clearly excluded.
@@
Figure 3: Scaling exponent $z$ extracted from the dynamic correlation functions for different values of the bare viscosity $\eta$ in model H0. We determined $z$ by comparing the correlation function for two different volumes, $L=40$ and $48$. We also show the prediction of the Kawasaki approximation, Eq. (17); the error band is defined by varying the correlation ength in the range $\xi\in[L/2\pi,L/2]$.
The values $z=3$ and $4$ are approximate, a more sophisticated calculation using the $\epsilon$-expansion gives $z=3.07$ and $z=3.96$, but the Kawasaki function is a very good approximation to the behavior of real fluids Swinney and Henry (1973). We conclude from Eq. (17) that in any finite volume we are likely to observe a scaling exponent between 3 and 4, and that observing the model H scaling exponent requires a combination of a very large correlation length and a small viscosity.
Since the renormalized viscosity is smaller in model H0 (see Fig. 1) we have explored the scaling behavior in model H0. At the critical point $m^{2}=m_{c}^{2}$ we compute the correlation for a range of values of $\eta$ and $L^{3}$. For any given $\eta$ we look for data collapse when comparing different $L$ in order to determine the value of $z$. This is shown in Fig. 2 for $\eta=0.01$ and $L=40$ and $L=48$. We observe that scaling works very well. We then plot the extracted value of $z$ as a function of $\eta$  [^4] The data in Fig. 3 are obtained by minimizing $|C(t,L\!=\!40)-C(\alpha t,L\!=\!48)|$ with respect to $\alpha=(40/48)^{z}$ in the regime $C(t)>0.15$ (with $C(0)\equiv 1$). The errorbar in $z$ is determined by propagating the errors in $C(t)$.. The result is shown in Fig. 3, and compared to the Kawasaki prediction. We observe the expected crossover from $z\simeq 4$ at large viscosity to $z\simeq 3$ for small viscosity. For $\eta=10^{-2}$ we obtain the dynamical exponent $z\simeq 3.013\pm 0.058$, consistent with the prediction of the two-loop $\epsilon$-expansion, $z\simeq 3.0712$ Adzhemyan et al. (1999).
@@
We consider model H with parameters relevant to a QCD critical endpoint. We take $a=0.75$ fm and $\Delta t=0.3$ fm, and consider a temperature $T=130$ MeV. Then, the enthalpy density of a non-interacting quark-gluon plasma corresponds to $\rho=11.1$ in lattice units. A viscosity to entropy density ratio $\eta/s=1/(4\pi)$ implies $\eta=0.50$ 555 Note that $\eta$ is somewhat bigger than the minimum value of $\eta_{R}$ shown in Fig. 1. This is consistent with the observation by Kovtun et al. Kovtun et al. (2011) that fluctuations lead to a bound on $\eta$ which is slightly weaker than the string theory bound $\eta/s\geq 1/(4\pi)$.. We take a simulation volume $V_{0}=L^{3}$ and measure the order parameter in half the simulation volume. The results are shown in Fig. 4. The correlation functions satisfy dynamical scaling for all values of $n$, but the relaxation time depends on $n$. In particular, $G_{2}(t)$ decays more quickly than $G_{1}(t)$, but the relaxation rate of $G_{3}(t)$ is intermediate between $G_{1}(t)$ and $G_{2}(t)$. These results are not compatible with simple mean field models. Note that at the critical point the correlation length is only limited by $L$, and equilibration is extremely slow: In physical units, the $1/e$ decay time in Fig. 4 exceeds $10^{3}$ fm.
@@ V Summary and outlook:
We have studied the time evolution of higher moments of the order parameter, but we have not attempted to perform a complete calculation that can be compared to higher order cumulants measured in relativistic heavy ion collisions. For this purpose we have to couple our calculation to a realistic background that describes an expanding and cooling fluid. There are two basic approaches that one might follow in order to achieve this. One approach is to extend the methods described in this work to stochastic relativistic fluid dynamics, where we retain all the degrees of freedom of the fluid (both shear and sound modes). This could be accomplished along the lines recently proposed in Basar et al. (2024). Another option is to couple the model described here to a deterministic background flow, obtained from conventional fluid dynamic simulations.
Acknowledgments: This work is supported by the U.S. Department of Energy, Office of Nuclear Physics through the Contracts DE-FG02-03ER41260 (T.S.) and DE-SC0020081 (V.S.). We used computing resources provided by the NC State University High Performance Computing Services Core Facility (RRID:SCR-022168), as well as resources funded by the Wesley O. Doggett endowment. We thank Andrew Petersen for assistance in working with the HPC infrastructure.